A girl weighing 407 N hangs suspended on the end of a rope 8 m if a friend pushes her laterally to one side until the rope makes an angle of 35o with the vertical, what lateral force is required at this point?


1
Expert's answer
2021-02-28T07:32:34-0500

Let's write the forces in projections on axis "x" and "y":


"F_{lateral}-Tsin\theta=0,""Tcos\theta-W_{girl}=0."

Let's first find the force of tension in the rope from the second equation:


"T=\dfrac{W_{girl}}{cos\theta}."

Substituting "T" into the first equation we can find the lateral force:


"F_{lateral}=W_{girl}tan\theta=407\ N\cdot tan35^{\circ}=285\ N."